# A Survey on Data-Centric Recommender Systems

### v1

@@ Abstract
Recommender systems (RS) have become essential tools for mitigating information overload in a range of real-world scenarios. Recent trends in RS have seen a paradigm shift, moving the spotlight from model-centric innovations to the importance of data quality and quantity. This evolution has given rise to the concept of data-centric recommender systems (Data-Centric RS), marking a significant development in the field. This survey provides the first systematic overview of Data-Centric RS, covering 1) the foundational concepts of recommendation data and Data-Centric RS; 2) three primary issues in recommendation data; 3) recent research developed to address these issues; and 4) several potential future directions in Data-Centric RS.
@@ 1 Introduction
Recommender systems (RS) have been widely adopted to alleviate information overload in various real-world applications, such as social media, e-commerce, and online advertising. The past few decades have witnessed the rapid development of recommendation models, evolving from traditional collaborative-filtering-based models Rendle et al. (2009) to more advanced deep-learning-based ones Ge et al. (2023), which markedly improved the accuracy, diversity, and interpretability of recommendation results Li et al. (2020).
However, as recommendation models are growing larger and increasingly complex – exemplified by P5 Geng et al. (2022), a recommendation model that integrates five recommendation-related tasks into a shared language generation framework – the primary constraint impacting recommendation performance gradually transitions towards the recommendation data. Rather than focusing solely on developing advanced models, an increasing number of researchers are advocating for the enhancement of recommendation data, leading to the emergence of the novel concept of data-centric recommender systems (Data-Centric RS) Zha et al. (2023).
The fundamental rationale behind Data-Centric RS is that data ultimately dictates the upper limits of model capabilities. Large and high-quality data constitutes the essential prerequisite for breakthroughs in performance. For instance, the remarkable advancements of GPT models in Natural Language Processing are mainly dependent on the use of huge and high-quality datasets Ouyang et al. (2022). Similarly in Computer Vision, Convolutional Neural Networks exhibit performance on par with Vision Transformers when they have access to web-scale datasets Smith et al. (2023). For RS, the implication is clear: the greater the quality Wang et al. (2021a) or the quantity Liu et al. (2023) of recommendation data, the more proficiently RS can characterize user preferences, resulting in recommendations that resonate well with users.
Despite considerable attention from researchers and a great variety of methods that have been put forth in Data-Centric RS, to the best of our knowledge, no effort has been made to gather and provide a summary of works in this promising and fast-developing research field. To fulfill this gap, we first conduct a systematic review of the literature on Data-Centric RS and write this survey. This structure is as follows. We first detail the specifics of data that could be used for recommendation and present the formalization of Data-Centric RS (Section 2). Next, we identify three key issues in recommendation data, i.e., data incompleteness, data noise, and data bias (Section 3), and categorize the existing literature in accordance with these issues (Section 4). Finally, we spotlight a number of encouraging research directions in Data-Centric RS (Section 5).
@@ 2.1 Recommendation Data
The goal of RS is to help users discover potential items of interest and then generate personalized recommendations. To achieve this goal, as shown in Figure 1, we summarize three types of data that can be used in RS:
  • User profiles: User profiles refer to a collection of information and characteristics that describe an individual user. In the context of recommendation, user profiles typically include personal details, preferences, behavior patterns, and other relevant data that help define and understand a user’s characteristics, needs, and interests.
  • Item attributes: Item attributes involve the specific details or characteristics that describe an item. These details can include color, material, brand, and other relevant information that helps identify or categorize the item.
@@
Figure 2: Model-Centric RS v.s. Data-Centric RS.
2.2 Data-Centric RS
As shown in Figure 2, different from Model-Centric RS which improve recommendation performance by refining models, Data-Centric RS shift the focus from model to data and emphasize the role of data enhancement. More specifically, given the recommendation data $\mathcal{D}$ (e.g., user-item interactions), Data-Centric RS aim to determine a strategy $f$, which takes the original data $\mathcal{D}$ as input and outputs the enhanced data $\mathcal{D}^{\prime}$:
@@ 2.2 Data-Centric RS
The enhanced data $\mathcal{D}^{\prime}$ could be used with different recommendation models to further improve their performance. We also attempt to answer the following questions to make readers further understand Data-Centric RS:
Q1: Are Data-Centric RS the same as Data-Driven RS? Data-Centric RS and Data-Driven RS are fundamentally different, as the latter only emphasize the use of recommendation data to guide the design of recommendation models without modifying the data, which are typically still Model-Centric RS Zha et al. (2023).
Q2: Why Data-Centric RS are necessary? The objective of the recommendation model is to fit the recommendation data. Without changing the data, Model-Centric RS may generalize or even amplify errors (e.g., bias) in the data, resulting in suboptimal recommendations Zhang et al. (2023a). Therefore, Data-Centric RS are necessary.
Q3: Will Data-Centric RS substitute Model-Centric RS? Data-Centric RS do not diminish the value of Model-Centric RS. Instead, these two paradigms are complementarily interwoven in building recommender systems. Model-centric methods can be used to achieve data-centric goals. For example, diffusion model Liu et al. (2023) can be an efficient tool for data augmentation. Data-centric methods can facilitate the improvement of model-centric objectives by enhancing the input data of models Lai et al. (2024).
Figure 3: Overview of data issues in RS.
@@ 3 Data Issues
As illustrated in Figure 3, we identify three key issues that recommendation data may suffer from, i.e., data incompleteness, data noise, and data bias.
@@ 3.1 Data Incompleteness
Data incompleteness issue refers to scenarios where data for making recommendations is inadequate, consequently resulting in information gaps or missing details regarding users or items. Specifically, the forms and reasons in which recommendation data can be incomplete include:
@@
  • Limited item attributes: In parallel, data regarding item attributes may also be lacking. This incompleteness hinders the ability to precisely portray items and capture their distinct characteristics Wu et al. (2020). For instance, an online bookshop may only provide basic data about a book such as the title and author, neglecting additional details like genre or publication year that can enhance the recommendation accuracy.
  • Sparse user-item interactions: RS can encounter the “cold start” problem when new users join. With limited or no historical interactions for these users, providing accurate recommendations becomes challenging Wang et al. (2019). Additionally, users usually do not engage with every item. For example, in a movie recommender system, a user may only rate a handful of the thousands of movies available, leading to an incomplete picture of true preferences.
  • Inadequate contextual information: Contextual information like ratings, location, or user reviews significantly contribute to generating effective recommendations. However, due to privacy issues or constraints in user feedback, the available contextual information may be incomplete or lack details Chae et al. (2019). For instance, a user might give a high rating for a hotel visit but does not provide a review that could offer useful insights about his/her preferences for future recommendations.
@@ 3.2 Data Noise
Data noise issue arises when a portion of the recommendation data is irrelevant or erroneous, which negatively impacts the quality of recommendations provided. The noisy recommendation data can appear in several forms:
  • Redundant data: An abundance of identical data is often the consequence of errors during the data collection process. RS might incorrectly identify and log certain identical item attributes multiple times, such as the same item category appearing repeatedly. Alternatively, it might record a user interacting with the same item multiple times – a shop page refresh mistake, for example.
@@
  • Spam or fake data: At a more malicious level, RS can be susceptible to spam or fake data, generated by either malicious users or automated bots trying to harm the system. This undesired data can significantly contaminate the pool of authentic user data and can drastically impact the quality of the recommendations, resulting in decreased user satisfaction Wu et al. (2023). A typical example is “review bombing”, where orchestrated attempts by certain users or bots fill a product’s review section with negative feedback to harm its overall ratings, even though the product may be generally well-received by the majority.
Figure 4: Illustration of data bias in RS.
@@ 4 Research Progress
We organize the existing literature according to the data issues in RS outlined before. Specific categorization as well as representative works and techniques are shown in Table 1.
@@ 4.1.1 Attribute Completion
Let $\mathcal{V}$ be the set of users and items. In practice, the profiles or attributes of some users or items are not available. Therefore, the set $\mathcal{V}$ can be divided into two subsets, i.e., $\mathcal{V}^{+}$ and $\mathcal{V}^{-}$, which denote the attributed set and the no-attribute set, respectively. Let $\mathcal{A}=\{a_{v}\mid v\in\mathcal{V}^{+}\}$ denote the input attribute set, attribute completion aims to complete the attribute for each no-attribute user or item $v\in\mathcal{V}^{-}$:
@@
Existing works on attribute completion mainly rely on utilizing the topological structure of given graphs (e.g., user-item interaction, knowledge graphs, and social networks) Wu et al. (2020); You et al. (2020); Jin et al. (2021); Tu et al. (2022); Zhu et al. (2023); Guo et al. (2023). For instance, by modeling user-item interactions with respective user or item attributes into an attributed user-item bipartite graph, AGCN Wu et al. (2020) proposes an adaptive Graph Convolutional Network for joint item recommendation and attribute completion, which iteratively performs two steps: graph embedding learning with previously learned attribute values, and attribute update procedure to update the input of graph embedding learning. Moreover, given a knowledge graph, HGNN-AC Jin et al. (2021) leverages the topological relationship between nodes as guidance and completes attributes of no-attribute nodes by weighted aggregation of the attributes of linked attributed nodes. AutoAC Zhu et al. (2023) identifies that different attribute completion operations should be taken for different no-attribute nodes, and models the attribute completion problem as an automated search problem for the optimal completion operation of each no-attribute mode. Instead of focusing on attribute completion accuracy, FairAC Guo et al. (2023) pays attention to the unfairness issues caused by attributes and completes missing attributes fairly.
@@ 4.1.2 Interaction Augmentation
Figure 5: Categorization of data denoising methods in RS.
@@
Different from negative sampling, positive augmentation focuses on how to choose an interaction as like Wang et al. (2019, 2021b); Yang et al. (2021); Zhang et al. (2021a); Lai et al. (2022); Liu et al. (2023); Wei et al. (2023). For example, EGLN Yang et al. (2021) selects uninteracted items with higher prediction scores and labels them as positive to enrich users’ interactions. CASR Wang et al. (2021b) leverages counterfactual reasoning to generate user interaction sequences in the counterfactual world. MNR-GCF Lai et al. (2022) constructs heterogeneous information networks and fully exploits the contextual information therein to identify potential user-item interactions. Based on Generative Adversarial Nets (GANs), AugCF Wang et al. (2019) generates high-quality augmented interactions that mimic the distribution of original interactions. Inspired by the superior performance diffusion models in image generation, DiffuASR Liu et al. (2023) adapts diffusion models to user interaction sequences generation, and a sequential U-Net is designed to capture the sequence information and predict the added noise of generated interactions. Leveraging the capabilities of LLMs, LLMRec Wei et al. (2023) also seeks to identify both positive and negative interactions from a candidate set using well-designed prompts.
@@ 4.1.3 Discussion
While a variety of methods exist for addressing the incomplete data issue, the fact remains that no single method is capable of comprehensively addressing all scenarios involving missing data. Consequently, a considerable amount of time and effort must be devoted to identifying missing information and selecting enhancement strategies. Furthermore, evaluating the quality of enhanced data is not straightforward – improvements in RS might be misleadingly attributed to the simple expansion of data volume, which can cloud the actual effects of refinements in data quality.
@@ 4.2.1 Reweighting-based Denoising
The reweighting-based method aims to assign lower weights to the noisy data (or assign higher weights to the noise-less dataWang et al. (2021a); Gao et al. (2022); Zhou et al. (2022); Zhang et al. (2023d); Ge et al. (2023); Wang et al. (2023c). Wang et al. Wang et al. (2021a) experimentally observe that noisy interactions are harder to fit in the early training stages, and based on this observation, they regard the interactions with large loss values as noise and propose an adaptive denoising training strategy called R-CE, which assigns different weights for noisy interactions according to their loss values during training. SLED Zhang et al. (2023d) identifies and determines the reliability of interactions based on their related structural patterns learned on multiple large-scale recommendation datasets. FMLP-Rec Zhou et al. (2022) adopts learnable filters for denoising in sequential recommendation. The learnable filters perform a Fast Fourier Transform (FFT) to convert the input sequence into the frequency domain and filter out noise through an inverse FFT procedure. SGDL Gao et al. (2022) leverages the self-labeled memorized data as guidance to offer denoising signals without requiring any auxiliary information or defining any weighting functions. AutoDenoise Ge et al. (2023) adopts reinforcement learning to automatically and adaptively learn the most appropriate weight for each interaction.
@@ 4.2.2 Selection-based Denoising
Instead of assigning lower weights, the selection-based method directly removes the noisy data Tong et al. (2021); Wang et al. (2021a); Yuan et al. (2021); Zhang et al. (2022b); Lin et al. (2023a); Zhang et al. (2023c); Quan et al. (2023); Wang et al. (2023b); Lin et al. (2023b). For instance, different from R-CE, Wang et al. Wang et al. (2021a) propose another adaptive denoising training strategy called T-CE, which discards the interactions with large loss values. RAP Tong et al. (2021) formulates the denoising process as a Markov Decision Process and proposes learning a policy network to select the appropriate action (i.e., removing or keeping an interaction) to maximize long-term rewards. DSAN Yuan et al. (2021) suggests using the entmax function to automatically eliminate the weights of irrelevant interactions. HSD Zhang et al. (2022b) learns both user-level and sequence-level inconsistency signals to further identify inherent noisy interactions. DeCA Wang et al. (2023b) finds that different models tend to make more consistent agreement predictions for noise-free interactions, and utilizes predictions from different models as the denoising signal. GDMSR Quan et al. (2023) designs a self-correcting curriculum learning mechanism and an adaptive denoising strategy to alleviate noise in social networks. STEAM Lin et al. (2023b) further designs a corrector that can adaptively apply “keep”, “delete”, and “insert” operations to correct the interaction sequence.

### v4

@@ Abstract
Recommender systems (RSs) have become an essential tool for mitigating information overload in a range of real-world applications. Recent trends in RSs have revealed a major paradigm shift, moving the spotlight from model-centric innovations to data-centric efforts (e.g., improving data quality and quantity). This evolution has given rise to the concept of data-centric recommender systems (Data-Centric RSs), marking a significant development in the field. This survey provides the first systematic overview of Data-Centric RSs, covering 1) the foundational concepts of recommendation data and Data-Centric RSs; 2) three primary issues of recommendation data; 3) recent research developed to address these issues; and 4) several potential future directions of Data-Centric RSs.
@@ 1 Introduction
Recommender systems (RSs) have been widely adopted to alleviate information overload in various real-world applications, such as social media, e-commerce, and online advertising. The past few decades have witnessed the rapid development of recommendation models, evolving from traditional collaborative-filtering-based models Rendle et al. (2009) to more advanced deep-learning-based ones Ge et al. (2023), which have markedly improved the accuracy, diversity, and interpretability of recommendation results Li et al. (2020).
However, as recommendation models are growing larger and increasingly complex, as exemplified by the P5 recommendation model Geng et al. (2022) that integrates five recommendation-related tasks into a shared language generation framework, the primary constraint impacting recommendation performance gradually transitions towards recommendation data. Instead of focusing solely on developing even more advanced models, an increasing number of researchers have been advocating for the enhancement of recommendation data, leading to the emergence of the novel concept of data-centric recommender systems (Data-Centric RSs) Zha et al. (2023).
The fundamental rationale behind Data-Centric RSs is that data ultimately dictates the upper limits of model capabilities. Large and high-quality data constitutes the essential prerequisite for breakthroughs in performance. For instance, the remarkable advancements of GPT models in natural language processing are mainly originated from the use of huge and high-quality datasets Ouyang et al. (2022). Similarly, in computer vision, convolutional neural networks exhibit performance on par with vision transformers when they have access to web-scale datasets Smith et al. (2023). For RSs, the implication is clear: the greater the quality Wang et al. (2021a) and/or the quantity Liu et al. (2023) of recommendation data, the more proficiently RSs can characterize user preferences, resulting in recommendations that resonate well with users.
Despite considerable attention from researchers and a great variety of methods that have been put forth in Data-Centric RSs, to the best of our knowledge, there has been no effort to gather and provide a summary of works in this promising and fast-developing research field. To fulfill this gap, we conduct a systematic review of the literature on Data-Centric RSs and provide insights from the following four aspects. We first detail the specifics of data that could be used for recommendation and present the formalization of Data-Centric RSs (Section 2). Next, we identify three key issues in recommendation data, i.e., data incompleteness, data noise, and data bias (Section 3), and categorize the existing literature in accordance with these issues (Section 4). Finally, we spotlight a number of encouraging research directions in Data-Centric RSs (Section 5).
@@ 2.1 Recommendation Data
The goal of RSs is to help users discover potential items of interest and then generate personalized recommendations. To achieve this goal, as shown in Figure 1, we summarize three types of data that can be used in RSs:
  • User profiles: User profiles refer to a collection of information and characteristics that describe an individual user. In the context of recommendation, user profiles typically include demographics, preferences, behavior patterns, and other relevant data that helps define and understand a user’s characteristics, needs, and interests.
  • Item attributes: Item attributes involve the specific details or characteristics that describe an item. These details can include color, material, brand, price, and other relevant information that helps identify or categorize the item.
@@
Figure 2: Model-Centric RSs v.s. Data-Centric RSs.
2.2 Data-Centric RSs
As shown in Figure 2, different from Model-Centric RSs which improve recommendation performance by refining models, Data-Centric RSs shift the focus from model to data and emphasize the importance of data enhancement. More specifically, given the recommendation data $\mathcal{D}$ (e.g., user-item interactions), Data-Centric RSs aim to determine a strategy $f$, which takes the original data $\mathcal{D}$ as input and outputs the enhanced data $\mathcal{D}^{\prime}$:
@@ 2.2 Data-Centric RSs
The enhanced data $\mathcal{D}^{\prime}$ could be used by different recommendation models to further improve their performance. We also attempt to answer the following questions to clarify the definition of Data-Centric RSs:
Q1: Are Data-Centric RSs the same as Data-Driven RSs? Data-Centric RSs and Data-Driven RSs are fundamentally different, as the latter only emphasize the use of recommendation data to guide the design of recommendation models without modifying the data, which are essentially still Model-Centric RSs Zha et al. (2023).
Q2: Why Data-Centric RSs are necessary? The objective of a recommendation model is to fit the recommendation data. Without changing the data, Model-Centric RSs may generalize or even amplify errors (e.g., noise or bias) in the data, resulting in suboptimal recommendations Zhang et al. (2023a). Therefore, Data-Centric RSs are necessary.
Q3: Will Data-Centric RSs substitute Model-Centric RSs? Data-Centric RSs do not diminish the value of Model-Centric RSs. Instead, these two paradigms complement each other to build more effective recommender systems. Interestingly, model-centric methods can be used to achieve data-centric goals. For example, diffusion models Liu et al. (2023) can be an effective tool for data augmentation. Data-centric methods can facilitate the improvement of model-centric outcomes. For instance, high-quality synthesized data could inspire further advancements in model design Sachdeva et al. (2022).
Figure 3: Overview of data issues in RSs.
@@ 3 Data Issues
As illustrated in Figure 3, we identify three key issues from which recommendation data may suffer, including data incompleteness, data noise, and data bias.
@@ 3.1 Data Incompleteness
The data incompleteness issue refers to the scenarios where data used for making recommendations is inadequate, consequently resulting in information gaps or missing details regarding users or items. Specifically, the forms and reasons in which recommendation data can be incomplete include:
@@
  • Limited item attributes: Similarly, data regarding item attributes may also be lacking. This incompleteness hinders the ability to precisely portray items and capture their distinct characteristics Wu et al. (2020). For instance, an online bookshop may only provide basic data about a book such as the title and author, neglecting additional details like genre or publication year that can enhance the recommendation accuracy.
  • Sparse user-item interactions: RSs can encounter the “cold start” problem when new users join. With limited or no historical interactions for these users, providing accurate recommendations becomes challenging Wang et al. (2019). Additionally, users usually do not engage with every item. For example, in a movie recommender system, a user may only rate a handful of thousands of movies available, leading to an incomplete picture of true preferences.
  • Inadequate contextual information: Contextual information like timestamps, ratings, locations, or user reviews significantly contributes to generating effective recommendations. However, due to privacy issues or constraints in user feedback, the available contextual information may be incomplete or lack details Chae et al. (2019). For instance, a user might give a high rating for a hotel visit but does not provide a review that could offer useful insights about his/her preferences for future recommendations.
@@ 3.2 Data Noise
Data noise arises when a portion of the recommendation data is irrelevant or erroneous, which negatively impacts the quality of recommendations provided. Noisy recommendation data can appear in several forms:
  • Redundant data: An abundance of identical data is often the consequence of errors during the data collection process. RSs might incorrectly identify and log certain identical item attributes multiple times, such as the same item category appearing repeatedly. Alternatively, it might record a user interacting with the same item multiple times – a shop page refresh mistake, for example.
@@
  • Spam or fake data: At a more malicious level, RSs can be susceptible to spam or fake data generated by either malicious users or automated bots trying to harm the system. This undesired data can significantly contaminate the pool of authentic user data and can drastically impact the quality of recommendations, leading to decreased user satisfaction Wu et al. (2023). A typical example is “review bombing”, where orchestrated attempts by certain users or bots fill a product’s review section with negative feedback to harm its overall rating, even though the product may be generally well-received by the majority.
Figure 4: An illustration of data bias in RSs.
@@ 4 Research Progress
We organize the existing literature according to the data issues in RSs outlined before. Specific categorization as well as representative works and techniques are shown in Table 1.
@@ 4.1.1 Attribute Completion
Let $\mathcal{V}$ be the set of users and items. In practice, the profiles or attributes of some users or items may not be available. Therefore, the set $\mathcal{V}$ can be divided into two subsets, i.e., $\mathcal{V}^{+}$ and $\mathcal{V}^{-}$, which denote the attributed set and the no-attribute set, respectively. Let $\mathcal{A}=\{a_{v}\mid v\in\mathcal{V}^{+}\}$ denote the input attribute set. Attribute completion aims to complete the attribute for each no-attribute user or item $v\in\mathcal{V}^{-}$:
@@
Existing works on attribute completion mainly rely on utilizing the topological structure of given graphs (e.g., user-item interaction, knowledge graphs, and social networks) Wu et al. (2020); You et al. (2020); Jin et al. (2021); Tu et al. (2022); Zhu et al. (2023); Guo et al. (2023). For instance, by modeling user-item interactions with respective user or item attributes into an attributed user-item bipartite graph, AGCN Wu et al. (2020) proposes an adaptive graph convolutional network for joint item recommendation and attribute completion, which iteratively performs two steps: graph embedding learning with previously learned attribute values, and attribute update procedure to update the input of graph embedding learning. Moreover, given a knowledge graph, HGNN-AC Jin et al. (2021) leverages the topological relationship between nodes as guidance and completes attributes of no-attribute nodes by weighted aggregation of the attributes of linked attributed nodes. AutoAC Zhu et al. (2023) identifies that different attribute completion operations should be taken for different no-attribute nodes and models the attribute completion problem as an automated search problem for the optimal completion operation of each no-attribute node. Instead of focusing on attribute completion accuracy, FairAC Guo et al. (2023) pays attention to the unfairness issue caused by attributes and completes missing attributes fairly.
@@ 4.1.2 Interaction Augmentation
Figure 5: Categorization of data denoising methods in RSs.
@@
Different from negative sampling, positive augmentation focuses on how to choose an interaction as like Wang et al. (2019, 2021b); Yang et al. (2021); Zhang et al. (2021a); Lai et al. (2022); Liu et al. (2023); Wei et al. (2023). For example, EGLN Yang et al. (2021) selects uninteracted items with higher prediction scores and labels them as positive to enrich users’ interactions. CASR Wang et al. (2021b) leverages counterfactual reasoning to generate user interaction sequences in the counterfactual world. MNR-GCF Lai et al. (2022) constructs heterogeneous information networks and fully exploits the contextual information therein to identify potential user-item interactions. Based on generative adversarial nets (GANs), AugCF Wang et al. (2019) generates high-quality augmented interactions that mimic the distribution of original interactions. Inspired by the superior performance of diffusion models in image generation, DiffuASR Liu et al. (2023) adapts diffusion models to user interaction sequence generation, and a sequential U-Net is designed to capture the sequence information and predict the added noise of generated interactions. Leveraging the capabilities of LLMs, LLMRec Wei et al. (2023) also seeks to identify both positive and negative interactions from a candidate set using well-designed prompts.
@@ 4.1.3 Discussion
While a variety of methods exist for addressing the incomplete data issue, the fact remains that no single method is capable of comprehensively addressing all scenarios involving missing data. Consequently, a considerable amount of time and effort must be devoted to identifying missing information and selecting enhancement strategies. Furthermore, evaluating the quality of enhanced data is not straightforward – improvements in RSs might be misleadingly attributed to the simple expansion of data volume, which can cloud the actual effects of refinements in data quality.
@@ 4.2.1 Reweighting-Based Denoising
The reweighting-based method aims to assign lower weights to the noisy data (or assign higher weights to the noiseless dataWang et al. (2021a); Gao et al. (2022); Zhou et al. (2022); Zhang et al. (2023d); Ge et al. (2023); Wang et al. (2023c). Wang et al. Wang et al. (2021a) experimentally observe that noisy interactions are harder to fit in the early training stages, and, based on this observation, they regard the interactions with large loss values as noise and propose an adaptive denoising training strategy called R-CE, which assigns different weights to noisy interactions according to their loss values during training. SLED Zhang et al. (2023d) identifies and determines the reliability of interactions based on their related structural patterns learned on multiple large-scale recommendation datasets. FMLP-Rec Zhou et al. (2022) adopts learnable filters for denoising in sequential recommendation. The learnable filters perform a fast Fourier transform (FFT) to convert the input sequence into the frequency domain and filter out noise through an inverse FFT procedure. SGDL Gao et al. (2022) leverages the self-labeled memorized data as guidance to offer denoising signals without requiring any auxiliary information or defining any weighting functions. AutoDenoise Ge et al. (2023) adopts reinforcement learning to automatically and adaptively learn the most appropriate weight for each interaction.
@@ 4.2.2 Selection-Based Denoising
Instead of assigning lower weights, the selection-based method directly removes the noisy data Tong et al. (2021); Wang et al. (2021a); Yuan et al. (2021); Zhang et al. (2022b); Lin et al. (2023a); Zhang et al. (2023c); Quan et al. (2023); Wang et al. (2023b); Lin et al. (2023b). For instance, different from R-CE, Wang et al. Wang et al. (2021a) propose another adaptive denoising training strategy called T-CE, which discards the interactions with large loss values. RAP Tong et al. (2021) formulates the denoising process as a Markov decision process and proposes to learn a policy network to select the appropriate action (i.e., removing or keeping an interaction) to maximize long-term rewards. DSAN Yuan et al. (2021) suggests using the entmax function to automatically eliminate the weights of irrelevant interactions. HSD Zhang et al. (2022b) learns both user-level and sequence-level inconsistency signals to further identify inherent noisy interactions. DeCA Wang et al. (2023b) finds that different models tend to make more consistent agreement predictions for noise-free interactions, and utilizes predictions from different models as the denoising signal. GDMSR Quan et al. (2023) designs a self-correcting curriculum learning mechanism and an adaptive denoising strategy to alleviate noise in social networks. STEAM Lin et al. (2023b) further designs a corrector that can adaptively apply “keep”, “delete”, and “insert” operations to correct an interaction sequence.